OpenStudy (keana):

1/5x + 1/2x + 2/5 = 16/5 A: the solution is x= ? B: the solution is all real numbers C: the equation has no solution

jimthompson5910 (jim_thompson5910):

Hint: The LCD is 10 (since the denominators are either 2 or 5, and 2*5 = 10). Multiply all of the terms by the LCD to clear out the fractions \[\Large \frac{1}{5}x+\frac{1}{2}x+\frac{2}{5} = \frac{16}{5}\] \[\Large {\color{red}{10*}}\frac{1}{5}x+{\color{red}{10*}}\frac{1}{2}x+{\color{red}{10*}}\frac{2}{5} = {\color{red}{10*}}\frac{16}{5}\] \[\Large 2x+5x+4 = 32\] I'll let you finish up

OpenStudy (anonymous):

a